Nicole has P100 and P500 bills. The number of P500 bills is 15 more than the number of P100 bills. If she has a total of P26 700 , how many pieces of bills does she have ?
Answer by ewatrrr(24785)

x number of P500 bills
500x + 100(x-15) = 26700
x = 28,200/600 = 47 P500 bills.
and...
32 P100 bills 47-15